OpenStudy (anonymous):

a spherical fish bowl has a radius of 21 cm.how many ml of water is needed to fill the half of the bowl?

OpenStudy (shinalcantara):

do you remember the formula for the volume of a sphere?

OpenStudy (shinalcantara):

V=(4/3)pi*r^3 is the volume for sphere V=(2/3)pi*r^3 for hemisphere now you substitute the given r to the volume formula for hemisphere and you'll get the volume

OpenStudy (shinalcantara):

for conversion: \[1 cm ^{3} = 1 ml\]